OpenStudy (anonymous):

Find a cubic function with the given zeros. -2, 3, -3 I actually found an answer for this but would like to make sure I solved it correctly My answer was f(x)=x^3+x-9. Please let me know if you get the same answer!

OpenStudy (amistre64):

subtract x from each root, and multiply the results

OpenStudy (amistre64):

-2, 3, -3 -2-x, 3-x, -3-x now when x is any of those values, we get at least one of them to be a zero; anything times zero is zero f(x) = (-2-x) (3-x) (-3-x)

OpenStudy (amistre64):

and any constant multiplier of it would work as well, i believe in this case that the "proper" outcome (a positive x^3 term) you might want to multiply it by -1 as well

OpenStudy (amistre64):

we could also test your results ... f(x)=x^3+x-9 ; let x = -2 (-2)^3+(-2)-9 ; let x = -2 -8-2-9 does not equal zero .... so it does not have -2 as a root